OpenStudy (anonymous):

Write an equation in point slope form of the line that passes through (2,-2) (5,7)

OpenStudy (zale101):



OpenStudy (anonymous):

the slope would be 3 I just don't know how to find the rest

OpenStudy (zale101):

where x1,y1=(2,-2) and x2,y2=(5,7)

OpenStudy (zale101):

this is the slope formula and m represents the slope

OpenStudy (zale101):

now plug int the numbers into that formula

OpenStudy (zale101):

Coordinates, not numbers. Sorry

OpenStudy (anonymous):

y-7=3(x-5)?

OpenStudy (rina.r):

(2,-2) (5,7) x1,y1 x2,y2 y2-y1/x2,x1 7-(-2) /5-2 =9 =3 so9/3 =3 is the answer.

OpenStudy (rina.r):

that's how u do it.

OpenStudy (anonymous):

so this is the answer y-7=3(x-5)

OpenStudy (rina.r):

no

OpenStudy (rina.r):

answer is 3 so do u want to write an equation of the line?

OpenStudy (anonymous):

Im confused because I know the slope is 3 I just don't know how to put it in point slope form

OpenStudy (rina.r):

y=mx+c

OpenStudy (rina.r):

y=3x+... sorry i forgot what to do after this..